You are to conduct an incident rate analysis of three companies for the year 2017. Assume all employees are full-time and there are no part-timers. Which company appears to have had the best safety program? Company AA had 9 total OSHA recordable cases with 235 employees. Company BB had 55 total OSHA recordable cases with 2900 employees. Company CC had 2 total OSHA recordable cases with 45 employees. Show your work.
Answer #1

OSHA Recordable Incident Rate (IR) = (Number of OSHA recordable cases*200000)/Number of employee labour hours worked

OSHA Recordable Incident Rate (IR) for company AA = (9*200000)/(235*50*40) = 3.829787234 [40 hours per week for 50 weeks per year]

OSHA Recordable Incident Rate (IR) for company BB = (55*200000)/(2900*50*40) = 1.896551724 [40 hours per week for 50 weeks per year]

OSHA Recordable Incident Rate (IR) for company CC = (2*200000)/(45*50*40) = 4.444444444 [40 hours per week for 50 weeks per year]

Company BB appears to have had the best safety program as it has lowest OSHA Recordable Incident Rate (IR)